# Learning Encodings by Maximizing State Distinguishability: Variational Quantum Error Correction

### v1

@@ 6 Hardware Experiments
In this section, we extend our experiments to actual quantum hardware to validate the practical applicability of our proposed VarQEC codes. We deploy our models on the 156-qubit ibm_marrakesh device featuring a Heron r2 chip provided by the IBM Quantum services [79]. The experiments involve running multiple logical patches in parallel on different regions of the chip. We compare the unencoded baseline to the logical patches, reconstructing the distinguishability loss using mitigated state tomography [48]. Details on the experimental procedure and the reconstruction method are summarized in Section F.1. The implementation of the deployment procedure on hardware and the raw results are provided as noted in the data availability statement.
@@
To control noise, we idle the qubits for a set duration, thereby inducing thermal relaxation noise. Mitigation techniques like dynamic decoupling were deactivated for the experiments. We identify the median properties of the ibm_marrakesh qubits to be $T_{1}=180\mu\text{s}$ and $T_{2}=120\mu\text{s}$. We adapt our ansatz directly to the heavy-hex lattice of the device (see also Section E.4), selecting CZ gates as the two-qubit gates and parameterized $\text{R}_{\text{zxz}}=R_{z}R_{x}R_{z}$ gates as single-qubit rotations to align with the native gate set. We train models on $n=3$, $n=4$, and $n=5$ qubits in simulation with a noise duration of $t=10.0\mu\text{s}$.
@@
In Fig. 8(a), we present results for several delay values ranging from $t=2.5\mu\text{s}$ to $t=10.0\mu\text{s}$ for a [[$4,1$]] VarQEC code trained at $t=10.0\mu\text{s}$. We observe that the baseline physical distinguishability loss increases sharply with increasing delay duration, while the distinguishability loss for encoded patches increases much more slowly, demonstrating the effectiveness of our encoding in preserving quantum information over time. Since our training in simulation assumes ideal encoding operations, we reduce the noise duration to $t=5.0\mu\text{s}$ in the subsequent experiment shown in Fig. 8(b) to partially account for additional sources of error such as noise introduced during the encoding operations themselves and measurement delays. We also train and deploy [[$3,1$]] and [[$5,1$]] VarQEC codes. The most important observation is that with an increasing number of physical qubits, the distinguishability loss decreases, indicating improved error correction performance with larger codes.
@@
Reliably predicting and quantifying scalability behavior from this limited data is challenging and is left for future work, as it requires extensive access to hardware platforms. In principle, these results could be improved by refining the noise model of the device, i.e. incorporating two-qubit correlated noise directly into the training procedure. We analyze the experiments and the results on the individual patches in more detail in Section F.2. Furthermore, we extend our experiments to a $20$-qubit IQM Resonance device [81] in Section F.3, where we demonstrate similar results. While the hardware limits qubit stability to shorter durations, its uniform qubit performance is beneficial for our procedure, since our training relies on consistent noise levels across all qubits. Overall, these experiments demonstrate the practicality and adaptability of our approach across multiple existing hardware platforms.
@@ 8 Summary and Conclusion
Data Availability Statement
The pipeline for training the proposed VarQEC codes, including the distinguishability loss function and all other required components, is available at https://github.com/nicomeyer96/varqec. The enclosed README provides setup details and instructions for reproducing the results of the paper. The scripts for employing trained codes on IBM and IQM hardware, as well as the raw results and analysis methods, are available separately at https://github.com/nicomeyer96/varqec-experiment. Further information and data is available upon reasonable request.
Acknowledgements
@@
Funding.
The research was supported by the Bavarian Ministry of Economic Affairs, Regional Development and Energy with funds from the Hightech Agenda Bayern via the project BayQS. The research is also part of the Munich Quantum Valley (MQV), which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
@@ Appendix A Background and Definitions
This appendix summarizes some important definitions and properties used throughout this paper that are already established in the literature. We start in Section A.1 by providing an overview of different measures on and between quantum states. The various quantum noise channels that are considered in this paper are formally introduced in Section A.2.
@@ A.1 Measures on Quantum States
where $\lambda_{i}$ are the eigenvalues of $\rho-\sigma$, and $r$ it the rank. The trace distance satisfies the properties of a metric, and is bounded by $0\leq T(\rho,\sigma)\leq 1$. A value of 00 thereby indicates that both states are identical, while a value of $1$ implies that the states are orthogonal [33]. The trace distance is invariant under unitary operations $U$, i.e. $T(U\rho U^{\dagger},U\sigma U^{\dagger})=T(\rho,\sigma)$. Furthermore, it is contractive for CPTP maps $\Phi$, such as quantum noise channels, following the data processing inequality [95], i.e. $T(\Phi(\rho),\Phi(\sigma))\leq T(\rho,\sigma)$.
@@
where $\lambda_{i}$ are the eigenvalues of the positive semi-definite matrix $\sqrt{\rho}\sigma\sqrt{\rho}$. For pure states $\rho=\ket{\psi}\bra{\psi}$ | and $\sigma=\ket{\phi}\bra{\phi}$ | it simplifies to
@@ A.2 Overview of Considered Noise Channels
$I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\leavevmode\nobreak\ X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\leavevmode\nobreak\ Y=\begin{bmatrix}0&-i\\ i&0\end{bmatrix},\leavevmode\nobreak\ Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$ (38)
The arguably simplest noise channel we consider, which is inspired by classical noise, is the bit-flip channel $\mathcal{N}_{\text{bit}}$. With a probability of $p$ a bit-flip, i.e. an $X$-error occurs, while with a probability of $1-p$ the state remains unchanged. On the Bloch sphere of a single qubit, this can be visualized as shrinking along the $y$ and $z$ axes. The associated Kraus matrices are as follows:
$K_{0}=\sqrt{1-p}\cdot I,\leavevmode\nobreak\ K_{1}=\sqrt{p}\cdot X$ (39)
@@
$K_{0}=\sqrt{1-p}\cdot I,\leavevmode\nobreak\ K_{1}=\sqrt{\frac{p}{3}}\cdot X,\leavevmode\nobreak\ K_{2}=\sqrt{\frac{p}{3}}\cdot Y,\leavevmode\nobreak\ K_{3}=\sqrt{\frac{p}{3}}\cdot Z$ (40)
@@
$K_{0}=\sqrt{1-p}\cdot I,\leavevmode\nobreak\ K_{1}=\sqrt{p_{x}}\cdot X,\leavevmode\nobreak\ K_{2}=\sqrt{p_{y}}\cdot Y,\leavevmode\nobreak\ K_{3}=\sqrt{p_{z}}\cdot Z$ (43)
Another noise type we consider is described by the amplitude damping channel $\mathcal{N}_{\text{ad}}$, which captures the irreversible loss of energy from a quantum system to its environment, leading to spontaneous transitions from high-energy to low-energy states. Concretely, the damping probability $0\leq\gamma\leq 1$ describes the likelihood of losing energy from the |$\ket{1}$ state towards the |$\ket{0}$ state, which itself remains stable under this noise model. This channel is non-unital, i.e. the center of the Bloch sphere is moved from its origin. The Kraus matrices are:
$K_{0}=\begin{bmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{bmatrix},\leavevmode\nobreak\ K_{1}=\begin{bmatrix}0&0\\ 0&\sqrt{\gamma}\end{bmatrix}$ (44)
The somewhat complementary phase damping channel $\mathcal{N}_{\text{pd}}$ describes the loss of quantum coherence without energy dissipation. It affects the relative phase between quantum states, but leaves the probability of being in |$\ket{0}$ and |$\ket{1}$ states unchanged. The strength of the noise is similarly captured by a damping probability $0\leq\gamma\leq 1$. It induces a shrinking of the Bloch sphere along the $x$ and $y$ axes, and is described by:
$K_{0}=\begin{bmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{bmatrix},\leavevmode\nobreak\ K_{1}=\begin{bmatrix}0&\sqrt{\gamma}\\ 0&0\end{bmatrix}$ (45)
@@
where typically $N_{j}$ is the same for all qubits. However, we additionally analyze the resilience of the encoding operations to multi-qubit correlated noise. As the employed VarQEC encoding ansätze consist of at most two-qubit interactions, we also restrict experiments to two-qubit correlated errors. In particular, we model a correlated two-qubit depolarizing channel, with strength $p$ and the following $164^{2}=$ 16 Kraus matrices:
$K_{0}=\sqrt{1-p}\cdot II,\leavevmode\nobreak\ K_{1-15}=\sqrt{\frac{1}{15}}\cdot P\leavevmode\nobreak\ \leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ P\in\{IX,IY,IZ,XI,\cdots,ZZ\}$ (47)
All noise channels represent CPTP maps, which implies that the trace distance underlies the data processing inequality as introduced in Section A.1, i.e.
@@ Appendix B Properties of Distinguishability Loss
In this appendix, we analyze the properties of the introduced distinguishability loss from a more theoretical perspective. An empirical evaluation with the loss integrated in the introduced VarQEC pipeline can be found in the main part and in Appendix E. For convenience, we re-state the definition of the distinguishability loss for a noise channel $\mathcal{N}$, both in the average-case formulation $\mathcal{D}$ and worst-case formulation $\overline{\mathcal{D}}$:
@@
$\overline{\mathcal{D}}(\mathcal{N}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5};\Theta})$ $=\max_{\rho,\sigma}\left(T(\rho,\sigma)-T(\mathcal{N}(\rho_{L}),\mathcal{N}(\sigma_{L}))\right)$ (50)
Hereby, $T$ denotes the trace distance, and $\rho_{L},\sigma_{L}$ are the encoded states corresponding to the pure initial states $\rho,\sigma$, respectively. These are produced by applying an encoding operation $U_{\mathrm{enc}}$ to the $k$-qubit state and additional $n-k$ ancilla qubits, initialized in state |$\ket{0}$:
$\rho_{L}=U_{\mathrm{enc}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Theta)}\left(\rho\otimes\ket{0}\bra{0}^{n-k}\right)U_{\mathrm{enc}}^{\dagger}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Theta)}$ (51)
Optionally, to allow for training, the encoding operation can depend on adjustable parameters $U_{\mathrm{enc}}(\Theta)$, e.g. variational rotation angles. To allow for a practically tractable evaluation, we define approximations based on a two-design $\mathcal{S}$ (see Section B.4 for details) which reads as:
@@
When using these loss functions as objectives for learning quantum error correction codes, which we dub as VarQEC codes, the actual target should be to optimize the worst-case behavior. Throughout this paper, we employ the average-case formulation as a proxy for this objective, as it allows for faster and more stable training. Both, the worst-case and average-case formulation trivially are lower bounded by 00 for the case that $T(\rho,\sigma)=T(\mathcal{N}(\rho),\mathcal{N}(\sigma))$ for all state pairs $\rho,\sigma$, i.e. there is no loss of information under noise channel $\mathcal{N}$.
In the following, we analyze different properties of this distinguishability loss. We start by analytically calculating the baseline for a single physical qubit under depolarizing noise in Section B.1. One of the central statements of this paper, the connection between the distinguishability loss and the existence of high-fidelity recovery operations, is derived in Section B.2. In Section B.3 we establish a direct connection between the value of the loss function and the distance of the respective code, and discuss the distinction between exact and approximate QEC. Finally, in Section B.4 we analyze the accuracy of the two-design approximations of the loss function.
@@ B.1 Baseline for Depolarizing Noise
Let us start with an equivalent re-formulation of the depolarizing channel in Eq. 40, which interprets the noise as driving towards a maximally mixed state:
@@
assuming that $p\leq\frac{3}{4}$. Plugging this into the definition of the lost trace distance from Eq. 14 yields
@@
This already allows us to derive the worst-case distinguishability loss for this setup. Using Eq. 15 and the fact that the trace distance takes its maximum value of $1$ for orthogonal pure states, we get
@@
where $\rho,\sigma$ are Haar-random states. For a noise strength of $p=0.1$, this yields $\overline{\mathcal{D}}(\mathcal{N}_{\text{dep}})\approx 0.133$, which coincides with the experimental predictions in Fig. 5(a).
@@
Let $T$ be the trace distance and let |$\ket{\psi},\ket{\phi}$ be Haar-random states in a Hilbert space of dimension $d$. The average of trace distances between pairs of these states depends on $d$ as
@@
$T(\ket{\psi},\ket{\phi})=\sqrt{1-\left|\expectationvalue{\psi|\phi}\right|^{2}}$ (66)
For independent Haar-random states, the squared overlap $F:=\left|\expectationvalue{\psi|\phi}\right|^{2}$ follows a Beta distribution $B(1,d-1)$ [109], i.e. we have the probability density function
@@
With the use of Eqs. 60 and 61 and Lemma 1 for the single-qubit case, i.e. a Hilbert space of dimension $d=2$, we can now continue:
[…]
@@
For a noise strength of $p=0.1$, we get an average-case distinguishability loss of $\mathcal{D}(\mathcal{N}_{\text{dep}})\approx 0.089$, which also matches the empirical result in Fig. 5(a).

### v2

@@ 6 Hardware Experiments
In this section, we extend our experiments to actual quantum hardware to validate the practical applicability of our proposed VarQEC codes. We deploy our models on the 156-qubit ibm_marrakesh device featuring a Heron r2 chip provided by the IBM Quantum services [85]. The experiments involve running multiple logical patches in parallel on different regions of the chip. We compare the unencoded baseline to the logical patches, reconstructing the distinguishability loss using mitigated state tomography [52]. Details on the experimental procedure and the reconstruction method are summarized in section˜F.1. The implementation of the deployment procedure on hardware and the raw results are provided as noted in the data availability statement.
(a) The $((4,2))$ VarQEC code is trained in simulation for a noise duration of $10.0\mu\text{s}$. We deploy the model on the QPU for delays varying from $2.5\mu\text{s}$ to $10.0\mu\text{s}$. The bar plots visualize the distribution of the worst-case distinguishability loss over multiple logical patches, with orange lines denoting the median and outliers exceeding $1.5$ times the interquartile range marked separately.
(b) VarQEC codes with increasing number of physical qubits, i.e. $((3,2))$, $((4,2))$, and $((5,2))$ are trained in simulation for a noise duration of $10.0\mu\text{s}$. To account for additional sources of errors, like noise introduced during the encoding operation itself and measurement delays, the delay is reduced to $5.0\mu\text{s}$ for hardware deployment. The violin plots visualize the distribution of the worst-case distinguishability loss over multiple logical patches.
Figure 8: Experiment on the ibm_marrakesh device [85], which features a 156-qubit Heron r2 chip. We induce thermal relaxation noise by applying a delay on all wires for a specific duration, assuming a median $T_{1}=180\mu\text{s}$ and $T_{2}=120\mu\text{s}$. The VarQEC codes are trained in classical simulation assuming these noise hyperparameters. Evaluation is conducted on hardware simultaneously on $8$ logical patches distributed uniformly over the QPU. We successively apply cluster re-sampling to enhance the granularity of data points [17, 73]. Additional details on the experimental procedure and the reconstruction of the worst-case distinguishability loss $\overline{\mathcal{D}}$ via mitigated state tomography can be found in the appendix. The predictions of the distinguishability loss from the simulation can be found in figure˜21(a).
@@
To control noise, we idle the qubits for a set duration, thereby inducing thermal relaxation noise. Mitigation techniques like dynamic decoupling were deactivated for the experiments. We identify the median properties of the ibm_marrakesh qubits to be $T_{1}=180\mu\text{s}$ and $T_{2}=120\mu\text{s}$. We adapt our ansatz directly to the heavy-hex lattice of the device, selecting CZ gates as the two-qubit gates and parameterized $\text{R}_{\text{zxz}}=R_{z}R_{x}R_{z}$ gates as single-qubit rotations to align with the native gate set. We train models on $n=3$, $n=4$, and $n=5$ qubits in simulation with a noise duration of $t=10.0\mu\text{s}$.
@@
In figure˜8(a), we present results for several delay values ranging from $t=2.5\mu\text{s}$ to $t=10.0\mu\text{s}$ for a $((4,2))$ VarQEC code trained at $t=10.0\mu\text{s}$. We observe that the baseline physical distinguishability loss increases sharply with increasing delay duration, while the distinguishability loss for encoded patches increases much more slowly, demonstrating the effectiveness of our encoding in preserving quantum information over time. Since our training in simulation assumes ideal encoding operations, we reduce the noise duration to $t=5.0\mu\text{s}$ in the subsequent experiment shown in figure˜8(b) to partially account for additional sources of error such as noise introduced during the encoding operations themselves and measurement delays. We also train and deploy $((3,2))$ and $((5,2))$ VarQEC codes. The most important observation is that with an increasing number of physical qubits, the distinguishability loss decreases, indicating improved error correction performance with larger codes. We did not explicitly report hardware results for the $[[5,1,3]]$, as after compilation to the restricted topology, the encoding circuit gets quite deep. Initial trials revealed that execution and successive tomography on hardware yield close-to-random measurement outcomes, due to accumulating noise.
@@
Reliably predicting and quantifying scalability behavior from this limited data is challenging and is left for future work, as it requires extensive access to hardware platforms. In principle, these results could be improved by refining the noise model of the device, i.e. incorporating two-qubit correlated noise directly into the training procedure. We analyze the experiments and the results on the individual patches in more detail in section˜F.2. Furthermore, we extend our experiments to a $20$-qubit IQM Resonance device [87] in section˜F.3, where we demonstrate similar results. While the hardware limits qubit stability to shorter durations, its uniform qubit performance is beneficial for our procedure, since our training relies on consistent noise levels across all qubits. Overall, these experiments demonstrate the practicality and adaptability of our approach across multiple existing hardware platforms.
@@ 8 Summary and Conclusion
[…]
@@
Some results of this study were previously reported in condensed form in [70]; here we present the complete, significantly extended version.
[…]
The research was supported by the Bavarian Ministry of Economic Affairs, Regional Development and Energy with funds from the Hightech Agenda Bayern via the project BayQS, and partially by the German Federal Ministry of Research, Technology and Space, funding program Quantum Systems, via the project Q-GeneSys, grant number 13N17389. The research is also part of the Munich Quantum Valley (MQV), which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
[…]
The pipeline for training the proposed VarQEC codes, including the distinguishability loss function and all other required components, is available at https://github.com/nicomeyer96/varqec. The enclosed README provides setup details and instructions for reproducing the results of the paper. The scripts for employing trained codes on IBM and IQM hardware, as well as the raw results and analysis methods, are available separately at https://github.com/nicomeyer96/varqec-experiment. Further information and data are available upon reasonable request.
@@ Appendix A Background and Definitions
This appendix summarizes some important definitions and properties used throughout this paper that are already established in the literature. We start in section˜A.1 by providing an overview of different measures on and between quantum states. The various quantum noise channels that are considered in this paper are formally introduced in section˜A.2.
@@ A.1 Measures on Quantum States
where $\lambda_{i}$ are the eigenvalues of $\rho-\sigma$, and $r$ it the rank. The trace distance satisfies the properties of a metric, and is bounded by $0\leq T(\rho,\sigma)\leq 1$. A value of 0 thereby indicates that both states are identical, while a value of $1$ implies that the states are orthogonal [36]. The trace distance is invariant under unitary operations $U$, i.e. $T(U\rho U^{\dagger},U\sigma U^{\dagger})=T(\rho,\sigma)$. Furthermore, it is contractive for CPTP maps $\Phi$, such as quantum noise channels, following the data processing inequality [103], i.e. $T(\Phi(\rho),\Phi(\sigma))\leq T(\rho,\sigma)$.
@@
where $\lambda_{i}$ are the eigenvalues of the positive semi-definite matrix $\sqrt{\rho}\sigma\sqrt{\rho}$. For pure states $|\rho=\ket{\psi}\bra{\psi}$ and $|\sigma=\ket{\phi}\bra{\phi}$ it simplifies to
@@ A.2 Overview of Considered Noise Channels
$I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\penalty 10000\ X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\penalty 10000\ Y=\begin{bmatrix}0&-i\\ i&0\end{bmatrix},\penalty 10000\ Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$ (42)
The arguably simplest noise channel we consider, which is inspired by classical noise, is the bit-flip channel $\mathcal{N}_{\text{bit}}$. With a probability of $p$ a bit-flip, i.e. an $X$-error occurs, while with a probability of 1p1-p the state remains unchanged. On the Bloch sphere of a single qubit, this can be visualized as shrinking along the $y$ and $z$ axes. The associated Kraus matrices are as follows:
$K_{0}=\sqrt{1-p}\cdot I,\penalty 10000\ K_{1}=\sqrt{p}\cdot X$ (43)
@@
$K_{0}=\sqrt{1-p}\cdot I,\penalty 10000\ K_{1}=\sqrt{\frac{p}{3}}\cdot X,\penalty 10000\ K_{2}=\sqrt{\frac{p}{3}}\cdot Y,\penalty 10000\ K_{3}=\sqrt{\frac{p}{3}}\cdot Z$ (44)
@@
$K_{0}=\sqrt{1-p}\cdot I,\penalty 10000\ K_{1}=\sqrt{p_{X}}\cdot X,\penalty 10000\ K_{2}=\sqrt{p_{Y}}\cdot Y,\penalty 10000\ K_{3}=\sqrt{p_{Z}}\cdot Z$ (47)
[…]
$K_{0}=\begin{bmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{bmatrix},\penalty 10000\ K_{1}=\begin{bmatrix}0&0\\ 0&\sqrt{\gamma}\end{bmatrix}$ (48)
[…]
$K_{0}=\begin{bmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{bmatrix},\penalty 10000\ K_{1}=\begin{bmatrix}0&\sqrt{\gamma}\\ 0&0\end{bmatrix}$ (49)
@@
where typically $N_{j}$ is the same for all qubits. However, we additionally analyze the resilience of the encoding operations to multi-qubit correlated noise. As the employed VarQEC encoding ansätze consist of at most two-qubit interactions, we also restrict experiments to two-qubit correlated errors. In particular, we model a correlated two-qubit depolarizing channel, with strength $p$ and the following $4^{2}=16$ Kraus matrices:
$K_{0}=\sqrt{1-p}\cdot II,\penalty 10000\ K_{1-15}=\sqrt{\frac{1}{15}}\cdot P\penalty 10000\ \penalty 10000\ \text{with}\penalty 10000\ P\in\{IX,IY,IZ,XI,\cdots,ZZ\}$ (51)
All noise channels represent CPTP maps, which implies that the trace distance underlies the data processing inequality as introduced in section˜A.1, i.e.
@@ Appendix B Properties of Distinguishability Loss
In this appendix, we analyze the properties of the introduced distinguishability loss from a more theoretical perspective. An empirical evaluation with the loss integrated in the introduced VarQEC pipeline can be found in the main part and in Appendix˜E. For convenience, we re-state the definition of the distinguishability loss for a noise channel $\mathcal{N}$, both in the average-case formulation $\mathcal{D}$ and worst-case formulation $\overline{\mathcal{D}}$:
@@
[…]
Optionally, to allow for training, the encoding operation can depend on adjustable parameters $U_{\mathrm{enc}}(\Theta)$, e.g. variational rotation angles. To allow for a practically tractable evaluation, we define approximations based on a two-design $\mathcal{S}$ (see section˜B.4 for details) which reads as:
@@
When using these loss functions as objectives for learning quantum error correction codes, which we dub as VarQEC codes, the actual target should be to optimize the worst-case behavior. Throughout this paper, we employ the average-case formulation as a proxy for this objective, as it allows for faster and more stable training. Both, the worst-case and average-case formulation trivially are lower bounded by 0 for the case that $T(\rho,\sigma)=T(\mathcal{N}(\rho),\mathcal{N}(\sigma))$ for all state pairs $\rho,\sigma$, i.e. there is no loss of information under noise channel $\mathcal{N}$.
In the following, we analyze different properties of this distinguishability loss. We start by analytically calculating the baseline for a single physical qubit under depolarizing noise in section˜B.1. One of the central statements of this paper, the connection between the distinguishability loss and the existence of high-fidelity recovery operations, is derived in section˜B.2. In section˜B.3 we establish a direct connection between the value of the loss function and the distance of the respective code, and discuss the distinction between exact and approximate QEC. Finally, in section˜B.4 we analyze the accuracy of the two-design approximations of the loss function.
@@ B.1 Baseline for Depolarizing Noise
Let us start with an equivalent re-formulation of the depolarizing channel in equation˜44, which interprets the noise as driving towards a maximally mixed state:
@@
assuming that $p\leq\frac{3}{4}$. Plugging this into the definition of the lost trace distance from equation˜18 yields
@@
This already allows us to derive the worst-case distinguishability loss for this setup. Using equation˜19 and the fact that the trace distance takes its maximum value of $1$ for orthogonal pure states, we get
@@
where $\rho,\sigma$ are Haar-random states. For a noise strength of $p=0.1$, this yields $\overline{\mathcal{D}}(\mathcal{N}_{\text{dep}})\approx 0.133$, which coincides with the experimental predictions in figure˜5(a).
@@
[…]
@@
[…]
For independent Haar-random states, the squared overlap $F:=\left|\expectationvalue{\psi|\phi}\right|^{2}$ follows a Beta distribution B(1,d1)B(1,d-1) [119], i.e. we have the probability density function
@@
With the use of equations˜64 and 65 and Lemma˜1 for the single-qubit case, i.e. a Hilbert space of dimension $d=2$, we can now continue:
$\mathcal{D}(\Theta;\mathcal{N}_{\text{dep}})$ $=\int_{\rho}\int_{\sigma}\Delta_{T}(\rho,\sigma;\mathcal{N}_{\text{dep}})\,d\nu(\rho)\,d\nu(\sigma)$ (76)
$=\frac{4}{3}p\cdot\int_{\rho}\int_{\sigma}T(\rho,\sigma)\,d\nu(\rho)\,d\nu(\sigma)$ (77)
@@
For a noise strength of $p=0.1$, we get an average-case distinguishability loss of $\mathcal{D}(\mathcal{N}_{\text{dep}})\approx 0.089$, which also matches the empirical result in figure˜5(a).